Data on Argentina from 2016 to 2018 is displayed in the table. Values are in US...

  1. Data on Argentina from 2016 to 2018 is displayed in the table. Values are in US billion dollars. Fill in the missing information and answer the questions below

Nominal GDP

(in US billion dollars)

GDP Deflator

Real GDP

(in US billion dollars)

2015

82.55

720.4

2016

557.7

707.2

2017

642.7

88.35

  1. Calculate the GDP growth rate from 2015 to 2016
  2. Calculate the GDP growth rate from 2016 to 2017
  3. What is the inflation rate from 2016 to 2017?
  4. What will Argentina’s GDP be in 5 years?
  5. How long will it take for Argentina’s GDP to double?
Answer #1

First use the fact that real GDP = nominal GDP / GDP deflator and find the missing values

Nominal GDP 2015 = 720.4*82.55/100 = 594.69

GDP deflator 2016 = 557.7*100/707.20 = 78.86

Real GDP 2017 = 642.7*100/88.35 = 727.45

GDP growth rate from 2015 to 2016 = real GDP growth rate = (707.2 - 720.4)*100/720.4 = -1.83%

Calculate the GDP growth rate from 2016 to 2017 = (727.45 - 707.20)*100/707.20 = 2.85%

Inflation rate from 2016 to 2017 = Using GDP deflator = (78.86 - 82.55)*100/82.55 = -4.47

Average growth rate of GDP in two years is (727.45/720.4)^(1/2) - 1 = 0.488%

Accordingly, if we start from 2017 GDP, GDP in 2022 will be 727.45*(1 + 0.488%)^5 = $745.37

Use rule of 72 to find the number of years required for Argentina’s GDP to double

= 72/0.488% = 148 years